# GradSafe: Detecting Unsafe Prompts for LLMs via Safety-Critical Gradient Analysis

### v1

@@ Abstract
Large Language Models (LLMs) face threats from unsafe prompts. Existing methods for detecting unsafe prompts are primarily online moderation APIs or finetuned LLMs. These strategies, however, often require extensive and resource-intensive data collection and training processes. In this study, we propose GradSafe, which effectively detects unsafe prompts by scrutinizing the gradients of safety-critical parameters in LLMs. Our methodology is grounded in a pivotal observation: the gradients of an LLM’s loss for unsafe prompts paired with compliance response exhibit similar patterns on certain safety-critical parameters. In contrast, safe prompts lead to markedly different gradient patterns. Building on this observation, GradSafe analyzes the gradients from prompts (paired with compliance responses) to accurately detect unsafe prompts. We show that GradSafe, applied to Llama-2 without further training, outperforms Llama Guard—despite its extensive finetuning with a large dataset—in detecting unsafe prompts. This superior performance is consistent across both zero-shot and adaptation scenarios, as evidenced by our evaluations on the ToxicChat and XSTest. The source code is available at https://github.com/xyq7/GradSafe.
Refer to caption Figure 1: Comparison of existing LLM-based unsafe prompt detection and GradSafe: a) Zero-shot LLM detectors can be imprecise, such as overestimating safety risks; b) Finetuned LLMs demand extensive training on carefully curated datasets; c) GradSafe accurately detects unsafe prompts using safety-critical gradients, without the need for LLM finetuning. Example prompt from XSTest (Röttger et al., 2023).
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Touvron et al., 2023) have achieved significant advancements in various domains (Klang and Levy-Mendelovich, 2023; Kung et al., 2023; Jiao et al., 2023; Goyal et al., 2022; Zhang et al., 2023). LLMs have also been integrated into various applications, such as search engine (Microsoft, 2023a) and office applications (Microsoft, 2023b). Moreover, finetuning LLMs for customized usage becomes possible with API finetuning services [^1] https://platform.openai.com/finetune or open-source LLMs (Touvron et al., 2023).
@@
However, unsafe user prompts pose threats to the safety of LLMs. On one hand, unsafe user prompts can lead to the misuse of LLMs, potentially facilitating various illegal or undesired consequences (Europol, 2023; Xie et al., 2023). Despite LLMs typically undergoing alignments with human values (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022), they remain vulnerable to various attacks (Selvi, 2022; Yi et al., 2023; Liu et al., 2023a), as well as instances of exaggerated safety (Röttger et al., 2023), which can overestimate the safety risks associated with user prompts. On the other hand, for LLM customization services, if unsafe prompts in the training set are not detected and filtered, the model can be readily finetuned to exhibit unsafe behavior and comply with unsafe prompts (Qi et al., 2023).
@@
To mitigate the risk of misuse and malicious finetuning, it is imperative to devise methods for the precise detection of unsafe prompts. While many API tools, including the Perspective API and OpenAI’s Moderation API (Markov et al., 2023), offer capabilities for online content moderation, these tools are primarily designed to detect general toxicity content, making them less effective in identifying unsafe prompts (Lin et al., 2023). With extensive knowledge base and reasoning capabilities, LLMs can also function as zero-shot detectors. However, LLMs employed as zero-shot detectors often exhibit suboptimal performance, such as an overestimation of safety risks. Recently, finetuned LLMs Inan et al. (2023); Pi et al. (2024), such as Llama Guard (Inan et al., 2023), have been proposed and demonstrate enhanced performance in detection tasks. Nonetheless, the finetuning process for LLMs requires a meticulously curated dataset and extensive training, necessitating substantial resources.
@@
In this work, we introduce GradSafe, which eliminates the need for dataset collection and finetuning of LLMs. In contrast to existing detectors that analyze the textual features of a prompt and/or an LLM’s response for it, GradSafe leverages gradients of the safety-critical parameters in LLMs. A comparison of existing LLM-based detectors and GradSafe is shown in Figure 1. The foundation of GradSafe is a critical observation: the gradients of an LLM’s loss for unsafe prompts paired with compliance response such as ‘Sure’ exhibit similar patterns (large cosine similarity) on particular parameter slices, in contrast to the divergent patterns observed with safe prompts. We characterize these parameters as ‘safety-critical parameters’.
@@
Leveraging this insight, GradSafe first meticulously analyzes the gradients of few reference safe and unsafe prompts (e.g., 2 examples for each, independent from evaluation dataset) coupled with compliance responses ‘Sure’. We identify safety-critical parameters as parameter slices that exhibit large gradient cosine similarities among unsafe prompts and small ones between unsafe and safe prompts. The average unsafe gradients for these parameter slices are stored as unsafe gradient reference. During detection, GradSafe pairs a given prompt with the compliance response ‘Sure’, computes the gradients of the LLM’s loss for this pair with respect to the safety-critical parameters, and calculates the cosine similarities with the unsafe gradient reference. We then introduce two variants of detection. The first, GradSafe-Zero, is a zero-shot, threshold-based classification method using the average of the cosine similarities across all slices as the score. Prompts with a score exceeding a predefined threshold are classified as unsafe. Alternatively, for situations requiring domain-specific adjustments, we present GradSafe-Adapt. This variant utilizes available data to construct a straightforward logistic regression model that employs the extracted cosine similarities as features to further enhance performance on the target domain.
@@
  • We make an observation that the gradients generated by unsafe prompts coupled with compliance responses exhibit consistent patterns on safety-critical parameters.
  • We propose GradSafe-Zero and GradSafe-Adapt, designed to detect unsafe prompts without necessitating further finetuning on an LLM with safety-critical gradient analysis.
@@ 2.1 Threats of Unsafe Prompts to LLM
Unsafe prompts pose threats to LLMs from mainly two aspects. On one hand, unsafe prompts can be leveraged for LLM misuse. Despite the safety alignment of LLMs (Bai et al., 2022; Kasirzadeh and Gabriel, 2022), LLMs can still be prompted to output harmful content (Perez and Ribeiro, 2022; Askell et al., 2021; Ganguli et al., 2022; Bai et al., 2022). There are various types of attacks, including jailbreak attacks (Xie et al., 2023; Liu et al., 2023b; Shen et al., 2023a) and prompt injection attacks (Liu et al., 2023a; Greshake et al., 2023; Iqbal et al., 2023; Yi et al., 2023), which can break the alignment of LLMs and facilitate misuse. Therefore, detecting unsafe prompts can serve as a first line of defense to prevent such misuse for LLM, which can be incorporated into different online ChatBot and LLM-integrated applications (Mialon et al., 2023; Schick et al., 2023; Shen et al., 2023b).
On the other hand, recent studies (Qi et al., 2023; Yi et al., 2024) demonstrate that malicious finetuning can significantly compromise the safety alignment when exposed to even a small number of unsafe prompts. However, existing online finetuning services fail to effectively detect such unsafe prompts, consequently leaving them vulnerable (Qi et al., 2023). As a result, the detection of unsafe prompts can be integrated into these finetuning services to screen out potentially harmful training data provided by users, thereby safeguarding LLMs against malicious finetuning.
@@ 2.2 Unsafe Prompt Detection
Before the widespread adoption of LLMs, content moderation efforts were primarily focused on certain types of online social media information (Jigsaw, 2017; Kiela et al., 2021; Hada et al., 2021), such as those found on platforms like Twitter (Zampieri et al., 2019; Basile et al., 2019), and Reddit (Hada et al., 2021). Various online moderation APIs are developed, such as OpenAI Moderation API, Azure API, Perspective API, etc. These APIs are typically based on models trained with vast amounts of data. For example, OpenAI has introduced the OpenAI Moderation API (Markov et al., 2023), which is designed to detect undesired content through meticulous data collection, labeling, model training, and active learning processes.
More recently, an increasing body of work has begun to pay attention to the detection of unsafe prompts in LLMs. ToxicChat (Lin et al., 2023) is proposed as a novel benchmark for the detection of unsafe prompts in LLMs, focusing on real user queries instead of content derived from social media platforms, which contains various potential unsafe prompts in conversation, including challenging cases such as jailbreaks. XSTest (Röttger et al., 2023) is proposed with unsafe and safe prompts to examine whether LLM suffers from exaggerated safety, which mistakes safe user prompts as unsafe. Recently, Llama Guard (Inan et al., 2023) has been introduced as an open-source model performing input-output unsafety detection specifically for LLMs, achieved by finetuning the Llama-2 model with a meticulously collected dataset. Unlike existing methods, our approach does not depend on further finetuning of LLMs. Instead, we show that we can accurately detect unsafe prompts by analyzing the safety-critical gradients of existing LLMs.
@@ 3.2 Identifying Safety-Critical Parameters
Step I (Obtaining gradients from unsafe/safe prompt response pairs):  We require only a minimal amount of prompts to acquire safety-critical parameters. To maintain generality and independence from the distribution of evaluation dataset, we only use two safe and two unsafe prompts. These prompts in our experiments are detailed in Appendix A. We compute an LLM’s standard loss for a pair of prompt and response ‘Sure’; and then calculate the gradient of the loss with respect to the LLM’s parameters.
The overall number of gradients/parameters for LLMs is huge and thus hard to analyze. Inspired by dimensional dependence observed in linguistic competence-related parameters (Zhao et al., 2023), for each gradient matrix, we slice them both row-wise and column-wise, leading to a total 2,498,56024985602,498,5602 , 498 , 560 slices (1,138,68811386881,138,6881 , 138 , 688 columns and 1,359,87213598721,359,8721 , 359 , 872 rows) for Llama-2 7b. These slices serve as the basic element in this work to identify safety-critical parameters and calculate cosine similarity features.
Figure 3: Illustration of the three phases in cosine similarities gap based filtering, where the threshold is $11$.
Step II (Cosine similarities gap based filtering):  Our objective is to identify parameter slices exhibiting high similarity in gradients across unsafe prompts, while demonstrating low similarity between unsafe and safe prompts. We present the process in multiple phases, using $33$ slices as an example in Figure 3. In Phase I, we obtain the average of the gradient slices for all unsafe prompts, which serve as reference gradient slices for subsequent cosine similarity computations. In Phase II, we compute the slice-to-slice cosine similarities between the gradient slices of each unsafe/safe sample and the corresponding reference gradient slices. In Phase III, our aim is to identify parameter slices with the largest gradient similarity gaps between unsafe and safe prompts. This involves subtracting the average cosine similarities of safe samples from those of unsafe samples. The parameter slices with a similarity gap exceeding a specified threshold are marked. The percents of marked slices for Llama-2 7b with different gap thresholds are detailed in Table 1. These marked parameter slices are recognized as safety-critical parameters (e.g., the third slice in Figure 3), and the corresponding gradient slices from the reference gradient slices are stored as unsafe gradient references.
@@ 4.1.1 Dataset
  • ToxicChat (Lin et al., 2023): ToxicChat is a dataset that comprises 10,1661016610,16610 , 166 prompts annotated with toxicity, curated from user interactions. We only use the prompts (user input) in the dataset for the experiment. The dataset is half split into training and testing sets. We use the official test set of ToxicChat-1123 for evaluation. For the adaption experiment, we use the official train set.
  • XSTest (Röttger et al., 2023): XSTest is a test suite encompassing a collection of $250250$ safe prompts from $1010$ types, and $200200$ corresponding crafted unsafe prompts. No training set is provided. We use the official test set of XSTest-v2 for evaluation.
@@ 4.1.2 Evaluation Metrics
In our evaluation, we adopt the Area Under the Precision-Recall Curve (AUPRC) as the primary metric for comparison against baseline models that can generate probabilities following the prior work (Inan et al., 2023). Moreover, we supplement our analysis by reporting precision, recall, and F1 scores to ensure a comprehensive assessment of performance. Specific settings to get the predictions for metric calculation for each baseline and GradSafe are detailed in Section 4.1.3 and 4.1.4.
4.1.3 Baselines
We include baselines from three categories: online API tools (OpenAI Moderation API, Perspective API, and Azure AI Content Safety API), LLMs as Zero-shot detectors (GPT4, and Llama-2), and finetuned LLM as detectors (Llama Guard).
  • OpenAI Moderation API [^2] https://platform.openai.com/docs/guides/moderation/: The OpenAI Moderation API is an online moderation tool based on the GPT model trained on content moderation datasets. It provides probabilities for $1111$ categories of safety risks. Following Llama Guard’s approach, we determine the overall unsafe score as the maximum probability across all categories. When computing precision, recall, and F1 score, we utilize the provided overall binary prediction label.
  • Perspective API [^3] https://perspectiveapi.com/: The Perspective API utilizes machine learning algorithms to identify harmful content across six categories of safety risks. We determine the overall unsafe score using the maximum probability across all categories. When computing precision, recall, and F1 score, a prompt is predicted as unsafe if the overall unsafe score exceeds $0.50.5$.
  • GPT-4 (OpenAI, 2023): GPT-4 is the state-of-the-art LLM. We use zero-shot prompting as a detector with the following prompt: ‘You are a helpful assistant. Provide your safety assessment for this user prompt: {Prompt} You must only answer with ‘safe’ or ‘unsafe’.’ In the computation of precision, recall, and F1 score, a prompt is classified as unsafe if ‘unsafe’ exists in the response. We use the GPT-4 API gpt-4-1106-preview in the evaluation.
  • Llama-2 (Touvron et al., 2023): Llama-2 is the base model for GradSafe and is the state-of-the-art open-source LLM. We also use zero-shot prompting as a detector with the same prompt and classification as GPT4. We use Llama-2-7b-chat-hf in the evaluation.
  • Llama Guard (Inan et al., 2023): Llama Guard is finetuned on the Llama-2 7b model using approximately 10,0001000010,00010 , 000 collected prompts and responses to generate classifications of ‘safe’ and ‘unsafe’ responses. Consistent with the methodology outlined in the original paper, we utilize the probability of producing ‘unsafe’ as the overall unsafe score and its binary output as its prediction result.
4.1.4 Settings for GradSafe
In GradSafe, we use Llama-2 (Llama-2-7b-chat-hf) as the base model. When identifying the safety-critical parameters, we use the gap threshold $11$. Given a prompt to detect, we use the system prompt ‘You are a helpful assistant. Help me with the following query: {Prompt} and pair it with the response ‘Sure’ to calculate the gradients. For GradSafe-Zero, we use the threshold $0.250.25$ for detection when calculating precision, recall, and F1 score on both benchmarks.
4.2 Overall Results
ToxicChat
XSTest
OpenAI Moderation API
0.604
0.779
Perspective API
0.487
0.713
@@ 4.2 Overall Results
ToxicChat
XSTest
OpenAI Moderation API
0.815/0.145/0.246
0.878/0.430/0.577
Perspective API
0.614/0.148/0.238
0.835/0.330/0.473
Azure API
0.559/0.634/0.594
0.673/0.700/0.686
@@
Llama-2
0.241/0.822/0.373
0.509/0.990/0.672
Llama Guard
0.744/0.396/0.517
0.813/0.825/0.819
GradSafe-Zero 0.753/0.667/0.707
0.856/0.950/0.900

### v2

# GradSafe: Detecting Jailbreak Prompts for LLMs via Safety-Critical Gradient Analysis

@@ Abstract
Large Language Models (LLMs) face threats from jailbreak prompts. Existing methods for detecting jailbreak prompts are primarily online moderation APIs or finetuned LLMs. These strategies, however, often require extensive and resource-intensive data collection and training processes. In this study, we propose GradSafe, which effectively detects jailbreak prompts by scrutinizing the gradients of safety-critical parameters in LLMs. Our method is grounded in a pivotal observation: the gradients of an LLM’s loss for jailbreak prompts paired with compliance response exhibit similar patterns on certain safety-critical parameters. In contrast, safe prompts lead to different gradient patterns. Building on this observation, GradSafe analyzes the gradients from prompts (paired with compliance responses) to accurately detect jailbreak prompts. We show that GradSafe, applied to Llama-2 without further training, outperforms Llama Guard—despite its extensive finetuning with a large dataset—in detecting jailbreak prompts. This superior performance is consistent across both zero-shot and adaptation scenarios, as evidenced by our evaluations on ToxicChat and XSTest. The source code is available at https://github.com/xyq7/GradSafe. Refer to caption Figure 1: Comparison of existing LLM-based jailbreak prompt detection and GradSafe: a) Zero-shot LLM detectors can be imprecise, such as overestimating safety risks; b) Finetuned LLMs demand extensive training on carefully curated datasets; c) GradSafe accurately detects jailbreak prompts using safety-critical gradients, without the need for LLM finetuning. Example prompt from XSTest Röttger et al. (2023).
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); OpenAI (2023); Chowdhery et al. (2022); Touvron et al. (2023) have achieved significant advancements in various domains Klang and Levy-Mendelovich (2023); Kung et al. (2023); Jiao et al. (2023); Goyal et al. (2022); Zhang et al. (2023). LLMs have also been integrated into various applications, such as search engine Microsoft (2023b) and office applications Microsoft (2023a). Moreover, finetuning LLMs for customized usage becomes possible with API finetuning services [^1] https://platform.openai.com/finetune or open-source LLMs Touvron et al. (2023).
@@
However, jailbreak/unsafe prompts pose threats to the safety of LLMs. On one hand, jailbreak prompts can lead to misuse of LLMs, potentially facilitating various illegal or undesired consequences Zou et al. (2023); Xie et al. (2023). Despite LLMs typically undergoing alignments with human values Ouyang et al. (2022); Bai et al. (2022), they remain vulnerable to various attacks Zou et al. (2023); Xie et al. (2023); Yi et al. (2023); Liu et al. (2024), as well as instances of exaggerated safety Röttger et al. (2023), which can overestimate the safety risks associated with user prompts. On the other hand, for LLM customization services, if jailbreak prompts in the training set are not detected and filtered, the model can be readily finetuned to exhibit unsafe behavior and comply with jailbreak prompts Qi et al. (2023).
@@
To mitigate the risk of misuse and malicious finetuning, it is imperative to devise methods for the precise detection of jailbreak prompts. While many API tools, including the Perspective API and OpenAI’s Moderation API Markov et al. (2023), offer capabilities for online content moderation, these tools are primarily designed to detect general toxicity content, making them less effective in identifying jailbreak prompts Lin et al. (2023). With extensive knowledge base and reasoning capabilities, LLMs can also function as zero-shot detectors. However, LLMs employed as zero-shot detectors often exhibit suboptimal performance, such as an overestimation of safety risks. Recently, finetuned LLMs like Llama Guard Inan et al. (2023) have been proposed and demonstrate enhanced performance in detection tasks. Nonetheless, the finetuning process for LLMs requires a meticulously curated dataset and extensive training, necessitating substantial resources.
@@
In this work, we introduce GradSafe, which eliminates the need for dataset collection and finetuning of LLMs. In contrast to existing detectors that analyze the textual features of a prompt and/or an LLM’s response for it, GradSafe leverages gradients of the safety-critical parameters in LLMs. A comparison of existing LLM-based detectors and GradSafe is shown in Figure 1. The foundation of GradSafe is a critical observation: the gradients of an LLM’s loss for jailbreak prompts paired with compliance response such as ‘Sure’ exhibit similar patterns (large cosine similarity) on particular parameter slices, in contrast to the divergent patterns observed with safe prompts. We characterize these parameters as ‘safety-critical parameters’.
@@
Leveraging this insight, GradSafe first meticulously analyzes the gradients of few reference safe and jailbreak prompts (e.g., 2 examples for each, independent from evaluation dataset) coupled with compliance responses ‘Sure’. We identify safety-critical parameters as parameter slices that exhibit large gradient cosine similarities among jailbreak prompts and small ones between jailbreak and safe prompts. The average unsafe gradients for these parameter slices are stored as unsafe gradient reference. During detection, GradSafe pairs a given prompt with the compliance response ‘Sure’, computes the gradients of the LLM’s loss for this pair with respect to the safety-critical parameters, and calculates the cosine similarities with the unsafe gradient reference. We then introduce two variants of detection. The first, GradSafe-Zero, is a zero-shot, threshold-based classification method using the average of the cosine similarities across all slices as the score. Prompts with a score exceeding a predefined threshold are classified as unsafe. Alternatively, for situations requiring domain-specific adjustments, we present GradSafe-Adapt. This variant utilizes available data to construct a straightforward logistic regression model that employs the extracted cosine similarities as features to further enhance performance on the target domain.
@@
  • We make an observation that the gradients generated by jailbreak prompts coupled with compliance responses exhibit consistent patterns on safety-critical parameters.
  • We propose GradSafe-Zero and GradSafe-Adapt, designed to detect jailbreak prompts without necessitating further finetuning on an LLM with safety-critical gradient analysis.
@@ 2.1 Threats of Unsafe Prompts to LLM
Unsafe/Jailbreak [^2] We use unsafe and jailbreak interchangeably in this paper. prompts pose threats to LLMs from mainly two aspects. On one hand, unsafe prompts can be leveraged for LLM misuse. Despite the safety alignment of LLMs Ouyang et al. (2022); Bai et al. (2022), LLMs can still be prompted to output harmful content Zou et al. (2023); Xie et al. (2023); Liu et al. (2023). Therefore, detecting unsafe prompts can serve as a first line of defense to prevent such misuse for LLM, which can be incorporated into different online ChatBot and LLM-integrated applications.
On the other hand, recent studies Qi et al. (2023); Yi et al. (2024) demonstrate that malicious finetuning can significantly compromise the safety alignment when exposed to even a small number of unsafe prompts with compliance responses. However, existing online finetuning services fail to effectively detect such unsafe prompts, consequently leaving them vulnerable Qi et al. (2023). As a result, the detection of unsafe prompts can be integrated into these finetuning services to screen out potentially harmful training data provided by users, thereby safeguarding LLMs against malicious finetuning.
@@ 2.2 Unsafe Prompt Detection
Before the widespread adoption of LLMs, content moderation efforts were primarily focused on certain types of online social media information (Jigsaw, 2017; Kiela et al., 2021; Hada et al., 2021), such as those found on platforms like Twitter (Zampieri et al., 2019; Basile et al., 2019), and Reddit (Hada et al., 2021). Various online moderation APIs are developed, such as OpenAI Moderation API, Azure API, Perspective API, etc.. These APIs are typically based on models trained with vast amounts of data. For example, OpenAI has introduced the OpenAI Moderation API Markov et al. (2023), which is designed to detect undesired content through meticulous data collection, labeling, model training, and active learning processes.
More recently, an increasing body of work has begun to pay attention to the detection of unsafe prompts in LLMs. ToxicChat Lin et al. (2023) is proposed as a novel benchmark for the detection of unsafe prompts in LLMs, focusing on real user queries instead of content derived from social media platforms, which contains various potential unsafe prompts in conversation, including challenging cases such as jailbreaks. XSTest Röttger et al. (2023) is proposed with unsafe and safe prompts to examine whether LLM suffers from exaggerated safety, which mistakes safe user prompts as unsafe. Recently, Llama Guard Inan et al. (2023) has been introduced as an open-source model performing input-output unsafety detection specifically for LLMs, achieved by finetuning the Llama-2 model with a meticulously collected dataset. Unlike existing methods, our approach does not depend on further finetuning of LLMs. Instead, we show that we can accurately detect unsafe prompts by analyzing the safety-critical gradients of existing LLMs.
@@ 3.2 Identifying Safety-Critical Parameters
Step I (Obtaining gradients from unsafe/safe prompt response pairs):  We require only a minimal amount of reference prompts to acquire safety-critical parameters. To maintain generality and independence from the distribution of evaluation dataset, we only use two safe and two unsafe prompts. These reference prompts in our experiments are detailed in Appendix A.2. We compute an LLM’s standard loss for a pair of prompt and response ‘Sure’; and then calculate the gradient of the loss with respect to the LLM’s parameters.
The overall number of gradients/parameters for LLMs is huge and thus hard to analyze. Inspired by dimensional dependence observed in linguistic competence-related parameters Zhao et al. (2023), for each gradient matrix, we slice them both row-wise and column-wise, leading to a total $2,498,560$ slices ($1,138,688$ columns and $1,359,872$ rows) for Llama-2 7b. These slices serve as the basic element in this work to identify safety-critical parameters and calculate cosine similarity features.
Figure 3: Illustration of the three phases in cosine similarities gap based filtering, where the threshold is $1$.
Step II (Cosine similarities gap based filtering):  Our objective is to identify parameter slices exhibiting high similarity in gradients across unsafe prompts, while demonstrating low similarity between unsafe and safe prompts. We present the process in multiple phases, using $3$ slices as an example in Figure 3. In Phase I, we obtain the average of the gradient slices for all unsafe prompts, which serve as reference gradient slices for subsequent cosine similarity computations. In Phase II, we compute the slice-to-slice cosine similarities between the gradient slices of each unsafe/safe sample and the corresponding reference gradient slices. In Phase III, our aim is to identify parameter slices with the largest gradient similarity gaps between unsafe and safe prompts. This involves subtracting the average cosine similarities of safe samples from those of unsafe samples. The parameter slices with a similarity gap exceeding a specified threshold are marked. The percents of marked slices for Llama-2 7b with different gap thresholds are detailed in Table 1. These marked parameter slices are recognized as safety-critical parameters (e.g., the third slice in Figure 3), and the corresponding gradient slices from the reference gradient slices are stored as unsafe gradient references.
@@ 4.1.1 Dataset
  • ToxicChat (Lin et al., 2023): ToxicChat is a dataset that comprises $10,166$ prompts annotated with toxicity, curated from user interactions. The dataset is half split into training and testing set. We use the official test set of ToxicChat-1123 for evaluation. For the adaption experiment, we use the official train set.
  • XSTest Röttger et al. (2023): XSTest is a test suite encompassing a collection of $250$ safe prompts from $10$ types, and $200$ corresponding crafted unsafe prompts. No training set is provided. We use the official test set of XSTest-v2 for evaluation.
@@ 4.1.2 Evaluation Metrics
In our evaluation, we adopt the Area Under the Precision-Recall Curve (AUPRC) as the primary metric for comparison against baseline models that can generate probabilities following the prior work Inan et al. (2023). Moreover, we supplement our analysis by reporting precision, recall, and F1 scores to ensure a comprehensive assessment of performance. Specific settings to get the predictions for metric calculation for each baseline and GradSafe are detailed in Section 4.1.3 and 4.1.4.
ToxicChat XSTest
OpenAI Moderation API 0.604 0.779
Perspective API 0.487 0.713
@@
ToxicChat XSTest
OpenAI Moderation API 0.815/0.145/0.246 0.878/0.430/0.577
Perspective API 0.614/0.148/0.238 0.835/0.330/0.473
Azure API 0.559/0.634/0.594 0.673/0.700/0.686
@@
Llama-2 0.241/0.822/0.373 0.509/0.990/0.672
Llama Guard 0.744/0.396/0.517 0.813/0.825/0.819
GradSafe-Zero 0.753/0.667/0.707 0.856/0.950/0.900
@@
4.1.3 Baselines
We include baselines from three categories: online API tools (OpenAI Moderation API, Perspective API, and Azure AI Content Safety API), LLMs as Zero-shot detectors (GPT4 and Llama-2), and finetuned LLM as detectors (Llama Guard).
  • OpenAI Moderation API [^3] https://platform.openai.com/docs/guides/moderation/: The OpenAI Moderation API is an online moderation tool based on the GPT model trained on content moderation datasets. It provides probabilities for $11$ categories of safety risks. Following Llama Guard’s approach, we determine the overall unsafe score as the maximum probability across all categories. When computing precision, recall, and F1 score, we utilize the provided overall binary prediction label.
  • Perspective API [^4] https://perspectiveapi.com/: The Perspective API utilizes machine learning algorithms to identify harmful content across six categories of safety risks. We determine the overall unsafe score using the maximum probability across all categories. When computing precision, recall, and F1 score, a prompt is predicted as unsafe if the overall unsafe score exceeds $0.5$.
  • GPT-4 OpenAI (2023): GPT-4 is the state-of-the-art LLM. We use zero-shot prompting as a detector with the following prompt: ‘You are a helpful assistant. Provide your safety assessment for this user prompt: {Prompt} You must only answer with ‘safe’ or ‘unsafe’.’ In the computation of precision, recall, and F1 score, a prompt is classified as unsafe if ‘unsafe’ exists in the response. We use the GPT-4 API gpt-4-1106-preview in the evaluation.
  • Llama-2 Touvron et al. (2023): Llama-2 is the base model for GradSafe and is the state-of-the-art open-source LLM. We also use zero-shot prompting as a detector with the same prompt and classification as GPT4. We use Llama-2-7b-chat-hf in the evaluation.
  • Llama Guard Inan et al. (2023): Llama Guard is finetuned on the Llama-2 7b model using approximately $10,000$ collected prompts and responses to generate classifications of ‘safe’ and ‘unsafe’ responses. Consistent with the methodology outlined in the original paper, we utilize the probability of producing ‘unsafe’ as the overall unsafe score and its binary output as its prediction result.
4.1.4 Settings for GradSafe
In GradSafe, we use Llama-2 (Llama-2-7b-chat-hf) as the base model. When identifying the safety-critical parameters, we use the gap threshold $1$. Given a prompt to detect, we use the system prompt ‘You are a helpful assistant. Help me with the following query: {Prompt} and pair it with the response ‘Sure’ to calculate the gradients. For GradSafe-Zero, we use the threshold $0.25$ for detection when calculating precision, recall, and F1 score on both benchmarks.
4.2 Overall Results